A 0.500 meter length of wire with a cross-sectional area of 3.14 x 10-6 meters squared is found to have a resistance of 2.53 x 10-3 ohms. According to the resistivity chart, from what material is the wire made?
1
Expert's answer
2021-10-25T14:51:05-0400

"L = 0.5 \;m \\

A = 3.14 \times 10^{-6} \;m^2 \\

R = 2.53 \times 10^{-3} \; \Omega \\

R = \frac{ρL}{A}"

ρ = resistivity of material

"ρ = \frac{RA}{L} \\

= \frac{2.53 \times 10^{-3} \times 3.14 \times 10^{-6}}{0.5} \\

= 15.8884 \times 10^{-9} \; \Omega-m"

According to the resistivity chart resistivity of silver is "15.9 \times 10^{-9} \; \Omega-m"

So, silver was used for formation of wire.